P = 100 – Q. where P is the price per calculator in dollars and Q is the number of calculators purchased per month. If the price is UD$ 30, how much revenue will calculator makers get each month? Find the price elasticity of demand for calculators at this point. 


1
Expert's answer
2021-06-18T12:30:41-0400

"Soln,"

Demand curve "P=100-Q"

P= $30 -Price of the calculators.

Q= No of sales of calculators p.m.

"P=100-Q"

"\$30=100-Q"

"Q=70"

At optimal quantity demanded price of "\$30=70" calculators sold p.m.

"TR=PQ"

"TR=\$30\times70"

Revenue collected p.m."=\$2100"


ii. "PED=\frac{\%\Delta Q}{\%\Delta P}"


"ed=\frac{\Delta Q}{\Delta P}\times\frac{P}{Q}"

"\frac {dQ}{dP}=\frac {d}{dQ}(100-Q)"

"=0-1"

"ed=-1\times\frac{30}{70}"

"=-0.43"